# MGDA Converges under Generalized Smoothness, Provably

### v4

@@ 5 Convergence Analysis under Iteration-wise CA distance
In Section 4 we show that the average CA distance is bounded under generalized smooth conditions. The average CA distance is also studied in MoCo [13] and MoDo [6] with the bounded gradient assumption and these works only focus on guarantees of the average CA distance over iterations. However, a $\epsilon$-level average CA distance only implies the smallest CA distance to be $\epsilon$-level. Since we want to keep the update direction close enough to the CA direction, it is better to have a tighter bound of CA distances. In this section, we show the iterative CA distance is $\mathcal{O}(\epsilon)$ with a warm-start process and convergence results for Algorithms 1, 3, and 4.
@@ 5.1 Deterministic setting
Let Assumptions 1 and 2 hold. Set $\beta^{\prime}\leq\frac{1}{M^{2}},\rho=\mathcal{O}(\epsilon^{2}),\beta\leq\mathcal{O}(\epsilon^{4}),\alpha\leq\mathcal{O}(\epsilon^{9})$, $N\geq O(\epsilon^{-2})$ as constants, and $T\geq\Theta(\epsilon^{-11})$. All the parameters satisfy the requirements in the formal version of Theorem 1 and we have $\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|\leq\mathcal{O}(\epsilon).$
The finite time error bound and the full proof can be found in Appendix D.1. Since our parameters satisfy the requirements in the formal version of Theorem 1, we can find an $\epsilon$-accurate Pareto stationary point with $\mathcal{O}(\epsilon^{-11})$ samples. In the analysis of CA distance, we show that the CA distance can be bounded by the term $\|w_{t}-w_{t,\rho}^{*}\|$ plus the strongly-convex constant $\rho$. Meanwhile, there is a decay relation between $\|w_{t+1}-w_{t+1,\rho}^{*}\|$ and $\|w_{t}-w_{t,\rho}^{*}\|$ with some error terms controlled by step sizes. Nevertheless, the error terms will accumulate since we do telescoping on this decay relation, which will be the dominating term. Thus, step sizes have to be much smaller than the choices in Theorem 1 to guarantee iteration-wise small CA distance.
@@
Let Assumptions 1 and 2 hold. Set $N\geq O(\epsilon^{-2})$, $\beta^{\prime}\leq\frac{1}{M^{2}},\rho=\min\left(\mathcal{O}(\epsilon^{2}),\mathcal{O}\left(\frac{1}{\alpha T}\right)\right),\beta\leq\mathcal{O}(\epsilon^{2}),\alpha\leq\min\left(\mathcal{O}(\beta),\mathcal{O}(\epsilon^{2}),\mathcal{O}\left(\frac{1}{\beta T}\right)\right)$ as constants, $T\geq\max\left(\Theta\left(\frac{1}{\alpha\epsilon^{2}}\right),\Theta\left(\frac{1}{\beta\epsilon^{2}}\right)\right)$. We have $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2}.$
The full version with detailed constants and proof can be found in the Section D.2. We can easily extend the analysis in Section C.1 to the convergence analysis of Algorithm 4 under the average CA distance, because the only extra effort is dealing with the remainder term, which can be simply bounded by the smallest step size. As a result, the sample complexity remains the same $\mathcal{O}(\epsilon^{-11})$ to achieve a Pareto stationary point.
@@
Let Assumptions 1 and 2 hold. We choose $\beta^{\prime}\leq\frac{1}{M^{2}},\rho=\mathcal{O}(\epsilon^{2}),\beta\leq\mathcal{O}(\epsilon^{4}),\alpha\leq\mathcal{O}(\epsilon^{9})$, $N\sim O(\epsilon^{-2})$ as constants, and $T\geq\Theta(\epsilon^{-11})$. We have that $\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|\leq\mathcal{O}(\epsilon).$
@@ 5.2 Stochastic setting
Let Assumptions 1, 2 and 3 hold. Set $\beta^{\prime}\leq\frac{1}{M^{2}}$, $\alpha\leq\mathcal{O}(\epsilon^{9})$, $\beta\leq\mathcal{O}(\epsilon^{4})$, $\rho=\mathcal{O}(\epsilon^{2})$, $n_{s}\geq\mathcal{O}(\epsilon^{-6}),N\geq\mathcal{O}(\epsilon^{-2}),$ and $T\geq\Theta(\epsilon^{-11})$, and all the parameters satisfy the requirements in Theorem 2. We then have $\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|\leq\mathcal{O}(\epsilon),$ with the probability at least $1-\delta$.
@@
The full version with detailed constants and proof can be found in Appendix D.4. Since our parameters satisfy all requirements in Theorem 2, we can find an $\epsilon$-accurate Pareto stationary point with high probability. Compared with Theorem 2, to guarantee an iteration-wise CA distance, despite our warm start process, a mini-batch method is required in our analysis. This is because given $\tau=T$, the gradient is not unbiased. In Theorem 2, the optimal stopping theorem is applied which indicates that the expectation of the cumulative gradient is zero. However, for each iteration, this optimal stopping theorem does not hold and the estimated error is controlled by the size of the mini-batch. Then, the sample complexity to get a Pareto stationary point becomes $\mathcal{O}(\epsilon^{-17})$ due to necessary mini-batch $n_{s}$.
@@
Method Segmentation Depth Surface Normal MR $\downarrow$ $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ Angle Distance $\downarrow$ Within $t^{\circ}$uparrow Mean Median 11.25 22.5 30 STL 38.30 63.76 0.6754 0.2780 25.01 19.21 30.14 57.20 69.15 LS 39.29 65.33 0.5493 0.2263 28.15 23.96 22.09 47.50 61.08 7.89 5.59 SI 38.45 64.27 0.5354 0.2201 27.60 23.37 22.53 48.57 62.32 7.33 4.39 RLW [23] 37.17 63.77 0.5759 0.2410 28.27 24.18 22.26 47.05 60.62 9.89 7.78 DWA [26] 39.11 65.31 0.5510 0.2285 27.61 23.18 24.17 50.18 62.39 7.11 3.57 UW [20] 36.87 63.17 0.5446 0.2260 27.04 22.61 23.54 49.05 63.65 7.11 4.05 MGDA [12] 30.47 59.90 0.6070 0.2555 24.88 19.45 29.18 56.88 69.36 5.56 1.38 MoCo [13] 40.30 66.07 0.5575 0.2135 26.67 21.83 25.61 51.78 64.85 5.00 0.16 MoDo [6] 35.28 62.62 0.5821 0.2405 25.65 20.33 28.04 54.86 67.37 7.55 0.49 Nash-MTL [29] 40.13 65.93 0.5261 0.2171 25.26 20.08 28.40 55.47 68.15 3.33 -4.04 FAMO [24] 38.88 64.90 0.5474 0.2194 25.06 19.57 29.21 56.61 68.98 3.22 -4.10 MGDA-warm start 40.57 67.17 0.5240 0.2281 25.21 19.74 28.74 55.79 68.21 2.78 -4.42
@@
6 Experiments
In this experiment, we evaluate the performance of the Cityscapes[10] and NYU-v2 [33] datasets. The former involves 2 pixel-wise tasks: 7-class semantic segmentation (Task 1) and depth estimation (Task 2) while the latter involves 3 pixel-wise tasks: 13-class semantic segmentation, depth estimation and surface normal estimation. Following the same experiment setup of [35], we build a SegNet [2] as the model. We compare the performance of MGDA with a warm start, which is used to ensure a small CA distance per iteration and mitigate gradient conflicts, against popular MGDA-type methods including MGDA [12], PCGrad [37], GradDrop [8], CAGrad [25], MoCo [13], MoDo [6], Nash-MTL [29], and FAMO [24]. Since SDMGrad [35] incorporates a preference-based regularization within the MGDA framework, we exclude it from the tables to ensure a fair comparison. We utilize the metric $\mathbf{\Delta m\%}$ to reflect the overall performance, which considers the average per-task performance drop versus the single-task (STL) baseline to assess methods. It can be observed in Table 2 and Figure 2 that MGDA-warm start has a much more balanced performance. Meanwhile, the proposed MGDA-FA is much faster as shown in Figure 4 in the Appendix.
We also illustrate the relationship between the gradient norm and the local smoothness for each task of the Cityscapes dataset. To do this, we compute them according to the method provided in Section H.3 in [38]. We scatter the local smoothness constant against gradient norms in Figure 1 for the semantic segmentation task and depth estimation task in Figure 3 (in the appendix), respectively. Both results demonstrate a positive correlation between them, which further substantiates the necessity of our analysis. More experimental details can be found in Appendix A.
Method Segmentation Depth $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ STL 74.01 93.16 0.0125 27.77 MGDA [12] 68.84 91.54 0.0309 33.50 44.14 PCGrad [37] 75.13 93.48 0.0154 42.07 18.29 GradDrop [8] 75.27 93.53 0.0157 47.54 23.73 CAGrad [25] 75.16 93.48 0.0141 37.60 11.64 MoCo [13] 75.42 93.55 0.0149 34.19 9.90 MoDo [6] 74.55 93.32 0.0159 41.51 18.89 Nash-MTL [29] 75.41 93.66 0.0129 35.02 6.82 FAMO [24] 74.54 93.29 0.0145 32.59 8.13 MGDA-warm start 75.41 93.46 0.0133 31.07 3.93
@@ 7 Conclusion
Appendix A Experimental details
A.1 Relation between gradient norms and the local smoothness
@@
A.2 Results and running time comparison between MGDA-warm start and MGDA-FA
We compare the results and average running time of the proposed algorithms, MGDA-warm start and MGDA-FA of the Cityscapes [10]. The time in Figure 4 is an average of the total running time over epochs (in minutes). The result solidifies the advantage of the fast approximation.
@@ A.2 Results and running time comparison between MGDA-warm start and MGDA-FA
A.3 Implementation details
Multi-task learning on Cityscapes dataset. Following the experiment setup in [35], we train our method for 200 epochs, using SGD optimizers for both model parameters and weights, and the batch size for Cityscapes is 8. We compute the averaged test performance over the last 10 epochs as the final performance measure. We fix the $\beta=0.5$ and do a grid search on hyperparameters including $N\in[10,20,40,50],\alpha\in[0.0001,0.0002,0.0005,0.001]$, and $\rho\in[0.01,0.05,0.1,0.2,0.5,0.6,0.7,0.8,0.9,1]$ and choose the best result from them. It turns out our best performance is based on the choice that $N=40,\alpha=0.0005,\beta=0.5$, and $\rho=0.5$. The choice of hyperparameters for MGDA-FA turns out to be the same as that for MGDA-warm start. All experiments are run on NVIDIA RTX A6000.

### v5

@@ 5 Convergence Analysis under Iteration-wise CA distance
In Section 4 we show that the average CA distance is bounded under generalized smooth conditions. The average CA distance is also studied in MoCo (Fernando et al., 2022) and MoDo (Chen et al., 2024) with the bounded gradient assumption and these works only focus on guarantees of the average CA distance over iterations. However, a $\epsilon$-level average CA distance only implies the smallest CA distance to be $\epsilon$-level. Since we want to keep the update direction close enough to the CA direction, it is better to have a tighter bound of CA distances. In this section, we show the iterative CA distance is $\mathcal{O}(\epsilon)$ with a warm-start process and convergence results for Algorithms 1, 3, and 4.
@@ 5.1 Deterministic setting
Let Assumptions 1 and 2 hold. Set $\beta^{\prime}\leq\frac{1}{M^{2}},\rho=\mathcal{O}(\epsilon^{2}),\beta=\mathcal{O}(\epsilon^{4}),\alpha=\mathcal{O}(\epsilon^{9})$, $N=\Omega(\epsilon^{-2})$ as constants, and $T=\Theta(\epsilon^{-11})$. All the parameters satisfy the requirements in the formal version of Theorem 1 and we have $\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|=\mathcal{O}(\epsilon).$
The finite time error bound and the full proof can be found in Appendix E.1. Since our parameters satisfy the requirements in the formal version of Theorem 1, we can find an $\epsilon$-accurate Pareto stationary point with $\mathcal{O}(\epsilon^{-11})$ samples. In the analysis of CA distance, we show that the CA distance can be bounded by the term $\|w_{t}-w_{t,\rho}^{*}\|$ plus the strongly-convex constant $\rho$. Meanwhile, there is a decay relation between $\|w_{t+1}-w_{t+1,\rho}^{*}\|$ and $\|w_{t}-w_{t,\rho}^{*}\|$ with some error terms controlled by step sizes. Nevertheless, the error terms will accumulate since we do telescoping on this decay relation, which will be the dominating term. Thus, step sizes have to be much smaller than the choices in Theorem 1 to guarantee iteration-wise small CA distance.
@@
Let Assumptions 1 and 2 hold. Set $N=\Omega(\epsilon^{-2})$, $\beta^{\prime}\leq\frac{1}{M^{2}},\rho=\mathcal{O}\big{(}\min\big{\{}\epsilon^{2},\frac{1}{\alpha T}\big{\}}\big{)},\beta=\mathcal{O}(\epsilon^{2}),\alpha=\mathcal{O}\big{(}\min\{\beta,\epsilon^{2},\frac{1}{\beta T}\})$ as constants, $T=\Theta\big{(}\max\{\frac{1}{\alpha\epsilon^{2}},\frac{1}{\beta\epsilon^{2}}\}\big{)}$. We have $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}=\mathcal{O}(\epsilon^{2}).$
The full version and proof can be found in Section E.2. We can easily extend the analysis in Section D.1 to the convergence analysis of Algorithm 4 under the average CA distance, because the only extra effort is dealing with the remainder term, which can be bounded by the smallest step size. As a result, the sample complexity remains the same $\mathcal{O}(\epsilon^{-11})$ to achieve a Pareto stationary point.
@@
Let Assumptions 1 and 2 hold. We choose $\beta^{\prime}\leq\frac{1}{M^{2}},\rho=\mathcal{O}(\epsilon^{2}),\beta=\mathcal{O}(\epsilon^{4}),\alpha=\mathcal{O}(\epsilon^{9})$, $N=\Omega(\epsilon^{-2})$ as constants, and $T=\Theta(\epsilon^{-11})$. We have that $\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|=\mathcal{O}(\epsilon).$
@@ 5.2 Stochastic setting
Let Assumptions 1, 2 and 3 hold. Set $\beta^{\prime}\leq\frac{1}{M^{2}}$, $\alpha=\mathcal{O}(\epsilon^{9})$, $\beta=\mathcal{O}(\epsilon^{4})$, $\rho=\mathcal{O}(\epsilon^{2})$, $n_{s}=\Omega(\epsilon^{-6}),N=\Omega(\epsilon^{-2}),$ and $T=\Theta(\epsilon^{-11})$, and all the parameters satisfy the requirements in Theorem 2. We then have $\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|=\mathcal{O}(\epsilon),$ with the probability at least $1-\delta$.
@@
The full version with detailed constants and proof can be found in Appendix E.4. Since our parameters satisfy all requirements in Theorem 2, we can find an $\epsilon$-accurate Pareto stationary point with high probability. Compared with Theorem 2, to guarantee an iteration-wise CA distance, despite our warm start process, a mini-batch method is required in our analysis. This is because given $\tau=T$, the gradient is not unbiased. In Theorem 2, the optional stopping theorem is applied which indicates that the expectation of the cumulative gradient is zero. However, for each iteration, this optional stopping theorem does not hold and the estimated error is controlled by the size of the mini-batch. Then, the sample complexity to get a Pareto stationary point becomes $\mathcal{O}(\epsilon^{-17})$ due to necessary mini-batch $n_{s}$.
6 Experiments
In this experiment, we evaluate the performance of the Cityscapes(Cordts et al., 2016) and NYU-v2 (Silberman et al., 2012) datasets. The former involves 2 pixel-wise tasks: 7-class semantic segmentation (Task 1) and depth estimation (Task 2) while the latter involves 3 pixel-wise tasks: 13-class semantic segmentation, depth estimation and surface normal estimation. Following the same experiment setup of (Xiao et al., 2024), we build a SegNet (Badrinarayanan et al., 2017) as the model. We compare the performance of MGDA with a warm start, which is used to ensure a small CA distance per iteration and mitigate gradient conflicts, against popular MGDA-type methods including MGDA (Désidéri, 2012), PCGrad (Yu et al., 2020), GradDrop (Chen et al., 2020), CAGrad (Liu et al., 2021), MoCo (Fernando et al., 2022), MoDo (Chen et al., 2024), Nash-MTL (Navon et al., 2022), and FAMO (Liu et al., 2024). Since SDMGrad (Xiao et al., 2024) incorporates a preference-based regularization within the MGDA framework, we exclude it from the tables to ensure a fair comparison. We utilize the metric $\mathbf{\Delta m\%}$ to reflect the overall performance, which considers the average per-task performance drop versus the single-task (STL) baseline to assess methods. It can be observed in Table 2 and Figure 2 that MGDA-warm start has a much more balanced performance. Meanwhile, the proposed MGDA-FA is much faster as shown in Figure 4 in the Appendix.
We also illustrate the relationship between the gradient norm and the local smoothness for each task of the Cityscapes dataset. To do this, we compute them according to the method provided in Section H.3 in Zhang et al. (2019). We scatter the local smoothness constant against gradient norms in Figure 1 for the semantic segmentation task and depth estimation task in Figure 3 (in the appendix), respectively. Both results demonstrate a positive correlation between them, which further substantiates the necessity of our analysis. More experimental details can be found in Appendix B.
@@ 6 Experiments
Method Segmentation Depth Surface Normal MR $\downarrow$ $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ Angle Distance $\downarrow$ Within $t^{\circ}$uparrow Mean Median 11.25 22.5 30 STL 38.30 63.76 0.6754 0.2780 25.01 19.21 30.14 57.20 69.15 LS 39.29 65.33 0.5493 0.2263 28.15 23.96 22.09 47.50 61.08 7.89 5.59 SI 38.45 64.27 0.5354 0.2201 27.60 23.37 22.53 48.57 62.32 7.33 4.39 RLW (Lin et al., 2021) 37.17 63.77 0.5759 0.2410 28.27 24.18 22.26 47.05 60.62 9.89 7.78 DWA (Liu et al., 2019) 39.11 65.31 0.5510 0.2285 27.61 23.18 24.17 50.18 62.39 7.11 3.57 UW (Kendall et al., 2018) 36.87 63.17 0.5446 0.2260 27.04 22.61 23.54 49.05 63.65 7.11 4.05 MGDA (Désidéri, 2012) 30.47 59.90 0.6070 0.2555 24.88 19.45 29.18 56.88 69.36 5.56 1.38 MoCo (Fernando et al., 2022) 40.30 66.07 0.5575 0.2135 26.67 21.83 25.61 51.78 64.85 5.00 0.16 MoDo (Chen et al., 2024) 35.28 62.62 0.5821 0.2405 25.65 20.33 28.04 54.86 67.37 7.55 0.49 Nash-MTL Navon et al. (2022) 40.13 65.93 0.5261 0.2171 25.26 20.08 28.40 55.47 68.15 3.33 -4.04 FAMO (Liu et al., 2024) 38.88 64.90 0.5474 0.2194 25.06 19.57 29.21 56.61 68.98 3.22 -4.10 MGDA-warm start 40.57 67.17 0.5240 0.2281 25.21 19.74 28.74 55.79 68.21 2.78 -4.42
@@
Method Segmentation Depth $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ STL 74.01 93.16 0.0125 27.77 MGDA (Désidéri, 2012) 68.84 91.54 0.0309 33.50 44.14 PCGrad (Yu et al., 2020) 75.13 93.48 0.0154 42.07 18.29 GradDrop (Chen et al., 2020) 75.27 93.53 0.0157 47.54 23.73 CAGrad (Liu et al., 2021) 75.16 93.48 0.0141 37.60 11.64 MoCo (Fernando et al., 2022) 75.42 93.55 0.0149 34.19 9.90 MoDo (Chen et al., 2024) 74.55 93.32 0.0159 41.51 18.89 Nash-MTL (Navon et al., 2022) 75.41 93.66 0.0129 35.02 6.82 FAMO (Liu et al., 2024) 74.54 93.29 0.0145 32.59 8.13 MGDA-warm start 75.41 93.46 0.0133 31.07 3.93$\pm$1.19
@@ 7 Conclusion
Acknowledgments
The work of Q. Zhang and S. Zou was partially supported by NSF under Grant CCF-2438429. P. Xiao and K. Ji were partially supported by NSF grants CCF-2311274 and ECCS-2326592.
Appendix A Notation summary
We have summarized the notations used in this paper in the following table.
Table 3: Notations and their descriptions.
Notations Descriptions
$x\in\mathbb{R}^{m}$ Model parameter, or decision variable
$s_{t,i,k}$ i-th collection of samples at time t in the stochastic model for training or testing
$f_{k}(x)$ A scalar-valued population objective function
$\nabla f_{k}(x)$ Gradient of $f_{k}(x)$, with $\nabla f_{k}(x):\mathbb{R}^{m}\mapsto\mathbb{R}^{m}$
$F(x)$ A vector-valued population objective function
$\nabla F(x)$ Gradient of $F(x)$, with $\nabla F(x):\mathbb{R}^{d}\mapsto\mathbb{R}^{d\times K}$
$\ell(a)$ A continuous non-decreasing function: [0,+)(0,+)00[0,$\infty$)$\rightarrow$(0,$\infty$)[ 0 , + ∞ ) → ( 0 , + ∞ )
$\varphi(a)$ A monotonically increasing for any $a\geq 0$: $\varphi(a)=\frac{a^{2}}{2\ell(2a)}$
$w\in\mathcal{W}$ Weighting parameter in a probability simplex over $[K]$,
$w^{*}_{\rho}\in\mathcal{W}$ optimal solution to equation 5, when $\rho=0$, it is simplified as $w^{*}$
$\alpha$ Step size to update model parameter $x$
$\beta$ Step size to update weight in the main loop
$\beta^{\prime}$ Step size to update weight in the warm start
$\rho$ Regularization parameter in equation 5
Appendix B Experimental details
B.1 Relation between gradient norms and the local smoothness
@@
B.2 Results and running time comparison between MGDA-warm start and MGDA-FA
We compare the results and average running time of the proposed algorithms, MGDA-warm start and MGDA-FA of the Cityscapes (Cordts et al., 2016). The time in Figure 4 is an average of the total running time over epochs (in minutes). The result solidifies the advantage of the fast approximation. For its performance, the fast approximation may introduce substantial errors in practice, potentially leading to inaccuracies in the weight update process. Meanwhile, the convergence analysis of the stochastic variant is not revealed. However, to enhance its effectiveness, we could apply a logarithmic technique in FAMO(Liu et al., 2024).
@@ B.2 Results and running time comparison between MGDA-warm start and MGDA-FA
B.3 Results of other MOO methods with warm start
The warm start strategy can be applied to other multi-objective optimization algorithms. For MGDA-based methods, both MoCo (Fernando et al., 2022) and MoDo (Chen et al., 2024) can incorporate the warm start as an add-on. However, the MoDo-warm start will exactly recover to our algorithm. Therefore, we evaluate the performance of the MoCo-warm start on the Cityscapes (Cordts et al., 2016) and NYU-v2 (Silberman et al., 2012) datasets. As shown in the following tables, the warm start strategy improves performance.
Method Segmentation Depth $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ MoCo 75.42 93.55 0.0149 34.19 9.90 MoCo-warm start 75.48 93.54 0.0148 31.43 7.32
Table 4: Multi-task supervised learning with MoCo-warm start on Cityscapes dataset.
Method Segmentation Depth Surface Normal $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ Angle Distance $\downarrow$ Within $t^{\circ}$uparrow Mean Median 11.25 22.5 30 MoCo 40.30 66.07 0.5575 0.2135 26.67 21.83 25.61 51.78 64.85 0.16 MoCo-warm start 38.40 64.40 0.5377 0.2315 26.04 20.57 27.11 54.02 66.63 -0.88
Table 5: Multi-task supervised learning with MoCo-warm start on NYU-v2 dataset.
B.4 Implementation details
Multi-task learning on Cityscapes dataset. Following the experiment setup in Xiao et al. (2024), we train our method for 200 epochs, using SGD optimizers for both model parameters and weights, and the batch size for Cityscapes is 8. We compute the averaged test performance over the last 10 epochs as the final performance measure. We fix the $\beta=0.5$ and do a grid search on hyperparameters including $N\in[10,20,40,50],\alpha\in[0.0001,0.0002,0.0005,0.001]$, and $\rho\in[0.01,0.05,0.1,0.2,0.5,0.6,0.7,0.8,0.9,1]$ and choose the best result from them. It turns out our best performance is based on the choice that $N=40,\alpha=0.0005,\beta=0.5$, and $\rho=0.5$. The choice of hyperparameters for MGDA-FA turns out to be the same as that for MGDA-warm start. All experiments are run on NVIDIA RTX A6000.
Multi-task learning on NYU-v2 dataset. Following the experiment setup in Xiao et al. (2024), we train our method for 200 epochs, using SGD optimizers for both model parameters and weights, and the batch size for NYU-v2 is 2. We compute the averaged test performance over the last 10 epochs as the final performance measure. We fix the $\beta=0.5$ and do a grid search on hyperparameters including $N\in[10,20,40,50],\alpha\in[0.0001,0.0002,0.0005,0.001]$, and $\rho\in[0.01,0.05,0.1,0.2,0.5,0.6,0.7,0.8,0.9,1]$ and choose the best result from them. It turns out our best performance is based on the choice that $N=40,\alpha=0.0001,\beta=0.5$, and $\rho=0.5$. All experiments are run on NVIDIA RTX A6000.